Fermi National Laboratory
 

Operation Summary for March 19, 2010 - March 22, 2010

 

The Integrated Luminosity for the period from 3/15/10 to 3/22/10 was 57.89 pb-1.
NuMI reported receiving 6.74E18 protons on target during this same period.

Friday March 19

The day shift began with Operations monitoring the Tevatron (TeV) store 7685 (68.78E30), the Recycler stash (418.7E10), and the Antiproton Source stack (6.0E10), with the Linac off due to problems with an RF station (LRF1), with Booster, Main Injector, NuMI, MiniBooNE, and the Meson MTest experiments waiting for beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, feeder 46A faulted somewhere between TeV sectors C0 and E0 just before 10:00 AM, resulting in half of the Mycom cryo system compressors around the TeV ring tripping off. Operations aborted store 7685 and bypassed the low beta power supplies. High voltage personnel isolated the fault and had power restored by 12:30 PM. CDF and D0 made collision hall accesses during the cryo recovery time. Linac personnel completed the switch tube repairs on a Linac RF station (LRF1) at 12:30 PM; this repair had begun on the previous shift. During the downtime NuMI personnel accessed MI-65 to grease the fan motor bearings. Operations resumed sending beam to Pbar, NuMI, and MiniBooNE. MI conducted studies during the afternoon. Pbar stacked and transferred antiprotons to the Recycler as appropriate. Operations sent beam to MTest, NuMI, and MiniBooNE when it was available.

During the evening shift, the TeV cryo system recovery continued. MI studies continued. Operations put the TeV into HEP shot setup at 7:38 PM. The last transfer of antiprotons was lost when the MI RF system tripped off in mid-transport; the store ended up being a 36x32 ( the first number is for bunches of protons and the second number is for bunches of antiprotons). Operations established store 7688 at 8:58 PM with an average initial luminosity (AIL) of 322.68E30. MI RF personnel came in and found a problem with an ADC card – they rebooted the low-level and mid-level RF systems and MI beam returned by 9:48 PM.

Saturday March 20

The midnight shift began with Operations monitoring store 7688 (176.36E30) and the stash (193.9E10), with Pbar stacking (8.8E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, the MIRF problem from the previous shift reared its ugly head again – Operations called MI RF personnel for advice and operators rebooted the low-level system again. Operations established beam to MTest around 4 AM.

During the day shift, Operations terminated store 7688 at 11:15 AM and put the TeV into shot setup. Unfortunately the first transfer of antiprotons to the TeV was lost due to MI RF problems; the TeV quenched. MI low-level RF personnel determined that new code had been causing the problems over they last few shifts, so they reverted to an older code set. The TeV recovered and Operations put the TeV into shot setup at 3:06 PM.

During the evening shift, Operations established store 7690 at 4:43 PM with an AIL of 345.68E30. Just before the end of the shift, ComEd lost a 345kV line somewhere off-site; this caused a power glitch that took out the store and tripped off systems throughout the complex. Operations established beam to users around 10:30 PM.

Sunday March 21

The midnight shift began with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations put the TeV into shot setup at 2:30 AM. Operations established store 7694 at 3:36 AM with an AIL of 253.2E30. Operations established beam to MTest around 4 AM.

During the day shift, operators rebooted the STATE front ends. Operations received a high water alarm for the downstream area of Enclosure E. The Duty Mechanic and Electrician investigated and reported the sump pump was working.

During the evening shift, the P1 transfer line beam valves closed for no apparent reason. The TeV A3 air compressor developed a leak at the drier. Operators valved in the bypass line and valved the drier out. Operations reported several minor Linac trips. Operators tuned the Booster to correct the beam positions in the MI-8 line. Operations terminated store 7694 at 8:05 PM. Operations put the TeV into shot setup and established store 7695 at 9:21 PM with an AIL of 350.7E30.

Monday March 22

The midnight shift began with Operations monitoring store 7695 and the stash, with Pbar stacking, with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations established beam to MTest at 5:54 AM. Operations reported minimal downtime due to various RF trips in various machines. Operations turned off the NuMI beam at 6 AM for a configuration change (from anti-neutrino to neutrino mode).

The Plans for Monday

The plans for today are stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/22/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory